OpenStudy (anonymous):

If the angles of a triangle are in the ratio 2:3:4. How many degrees are there in the largest angle?

OpenStudy (anonymous):

@ganeshie8 can you check my answer?

ganeshie8 (ganeshie8):

ha wat do u get ?

OpenStudy (anonymous):

See we'll add them first so I'm getting 9 then we'll divide with each one of them 2/9*180=40 3/9*180=60 4/9*180=80 So, 80 is the answer naa?

ganeshie8 (ganeshie8):

ha wahi.. good job !!

OpenStudy (anonymous):

Thanks for the help :)

ganeshie8 (ganeshie8):

np :)